# Multi-Excitation Projective Simulation with a Many-Body Physics-Inspired Inductive Bias

### v3

@@ V.3 Computer Maintenance
Figure 7 illustrates the process leading up to the agent receiving a reward from the environment. The reward is decomposed into two separate mechanisms that are each applied to different $h$-values, which we call the hypothesis and plausibility rewards. We do this to avoid washing out the percept information that tends to happen when blindly applying the standard PS update rule to intermediate layers. The hypothesis reward is measured based on how close each intermediate layer element is to its corresponding partner in the given scenario; it is applied to the $h$-values between the percept and intermediate layers. If those elements match exactly, a reward of +5 is given, otherwise, a large penalty of -10 is applied. An additional component of the reward penalizes the agent by up to -1 if they pick more elements than there are in the given scenario; an attempt to discourage the agent from selecting the maximum number of elements permitted by their inductive bias (becomes more important for increasing many-body cutoff size).
The plausibility reward, which is applied to the $h$-values between the intermediate layer and the action layer, has three different components to it. The first component has the same structure as the hypothesis reward but the values are quadrupled for the penalty on too many elements and it is for the agent’s chosen elements from the action layer instead of the intermediate layer. The second component checks whether the chosen components from the intermediate layer match those chosen from the action layer; a reward of +1 is given if they match perfectly, +0.25 if all of the action layer elements match but more intermediate layer elements are chosen then action layer elements, and -2 otherwise. This component of the reward can be interpreted as an internal consistency mechanism that encourages the agent to form coherent explanations between hypothesis and action. Note that it does not directly refer to the scenario key and is thus a general mechanism that can aid learning on scenarios not considered in the original training set. The third component of the reward checks whether the agent’s chosen fixes make sense with respect to the underlying causes they identified in their explanation. This is judged based on certain causal relationships between what the causes and fixes each refer to, that are put in by hand. For example, implementing the fix ‘replace components’ would be justified if the cause of the problem was suspected to be ‘physical damage’ or ‘faulty,’ since both indicate problems with the hardware that are only fixable by physically removing them. However, if the suspected cause was simply ‘malware,’ then only selecting ‘replace components’ would not be justified because this is a fix one implements when there are hardware problems, not software. Now, for each possible fix there are specific causes that need to be selected for the agent to receive the associated reward of +0.3 divided by the number of fixes specified in the chosen scenario key $n_{\mathrm{fix}}$; this ensures that the agent can converge to the optimal reward since some scenarios require different numbers of fixes to solve than others. A penalty of up to $-4/n_{\mathrm{fix}}$ is given if the proper causes are not selected. Aside from the use of $n_{\mathrm{fix}}$, the third component of the plausibility reward is also independent of the scenario key, adding another general mechanism for augmenting learning beyond the training set.
Finally, if all elements in the agent’s explanation exactly match those in the scenario key, the agent receives a big bonus to both rewards of +15 to amplify the probability of selecting this deliberation path again in future episodes. Once this bonus is triggered it signals an end to the episode and a fresh set of symptoms corresponding to another scenario are then selected. Before receiving this signal and before the agent factors in the reward into their update rule definition 6, the reward is first sent to an external reward shaping function defined by Equation (11) whose purpose is to discourage the agent from taking too many steps within an episode, intending to curtail arbitrarily long episodes; details about this function can be found in Appendix A.
All agents for each of their layers use the softmax function with $\beta=1/2$ to convert $h$-values to probabilities and all $h$-values are initialized to 1. The results of the training are shown in Figure 8. We can see that the inductive bias agent achieves near-optimal values for both the hypothesis and plausibility rewards using far fewer steps but requiring more episodes than the unrestricted agent does, as seen in the total step number per episode curve, which better illustrates the savings from the inductive bias as the two average reward plots do not show how many total steps have elapsed during training. The total step number per episode curve also shows roughly exponential decay, which is indicative of the agent first trying out scenario configurations randomly and then always picking the right configuration immediately. To quantify the difference in step number to the optimal reward, we calculated the total average steps taken over 200 episodes for both agents: the inductive bias agent takes roughly 5474 steps while the unrestricted agent takes roughly 8168 steps. Together with the complexity-theoretic results from Appendix D, we only expect this difference in step number to grow. Although the unrestricted agent does converge to the optimal rewards in fewer episodes, it clearly takes more steps to do so because it has roughly 16 times the number of trainable parameters as the inductive bias agent has. The fact that we get comparable performance from the inductive bias agent using far fewer parameters justifies the relatively small fluctuations around the optimal value for each of the reward curves, which are expected to decrease as more agents are included in the average.
@@
VI A Quantum Motivation of the Inductive Bias and a Path Towards Quantization
In Section IV, we gave a qualitative, conceptual motivation of our inductive biases. There, we identified that the size of the power set $\mathcal{P}(V)$ can lead to exponential costs, because in general all of its elements might be valid excitation configurations for which we have to define $h$-values.
However, the analogy to quantum many-body physics can be worked out more formally, and this in particular suggests a natural quantization for analog quantum computers. Therefore, in Section VI.1 we will work out this formal analogy, while in Section VI.2 we will provide some basic considerations for an actual quantum implementation.
VI.1 Full Quantum Motivation
In this section, we will provide more formal details on the analogy to quantum many-body physics which inspired our inductive biases. For that purpose, we start by explaining how excitation configurations are represented in quantum physics. Specifically, each atomic clip in our hypergraph can be interpreted as a mode in quantum many-body physics and in the formalism called second quantization [32, 33], each excitation configuration is represented by a vector |$\ket{n_{1},\dots,n_{|V|}}$ in a Hilbert space, where $n_{j}$ is the number of excitations in atomic clip/mode $j$.
Time evolution over the time duration $\Delta t$ is described by an operator $U(\Delta t)=e^{-i\Delta t\cdot H}$, where $i$ is the complex unit, $H$ is an operator called the Hamiltonian, and $e^{\bullet}$ is the matrix exponential. This means after a duration $\Delta t$, a many-body system starting in a state |$\ket{n_{1},\dots,n_{|V|}}$ is afterwards described by the state $e^{i\Delta t\cdot H}\ket{n_{1},\dots,n_{|V|}}$.
Important for our inductive bias are the typical expressions for $H$. For this, we require the ladder operators $a_{j}$ and $a^{\dagger}_{j}$, with denoting the hermitian adjoint (i.e. $A^{\dagger}=\overline{A^{T}}$, with $\overline{\bullet}$ denoting complex conjugation) and $j\in V$. $a^{\dagger}_{j}$ adds one excitation to atomic clip $j$, i.e. $a^{\dagger}_{j}\ket{\ldots,n_{j},\ldots}\propto\ket{\ldots,n_{j}+1,\ldots}$ and is called a creation operator. Similarly, $a_{j}$ removes an excitation from an atomic clip $j$, i.e. $a_{j}\ket{\ldots,n_{j},\ldots}\propto\ket{\ldots,n_{j}-1,\ldots}$, and is called an annihilation operator. For the special case $n_{j}=0$, we have $a_{j}\ket{\ldots,n_{j},\ldots}=0$.
@@ VI.1 Full Quantum Motivation
$H=\sum_{o,i}\sum_{j_{1},\ldots,j_{i}}\sum_{k_{1},\ldots,k_{o}}$ $h(j_{1},\ldots,j_{i},k_{1},\ldots,k_{o})$ (8)
×$\times a^{\dagger}_{k_{1}}\ldots a^{\dagger}_{k_{o}}\cdot a_{j_{1}}\ldots a_{j_{i}}$
@@
In many-body physics, given a state of the form $\sum_{n_{1},\ldots n_{|V|}}\alpha_{n_{1},\ldots,n_{|V|}}\ket{n_{1},\ldots,n_{|V|}}$ with $\alpha_{n_{1},\ldots,n_{|V|}}\in\mathbb{C}$ and $\sum_{n_{1},\ldots,n_{|V|}}|\alpha_{n_{1},\ldots,n_{|V|}}|^{2}=1$, each $|\alpha_{n_{1},\ldots,n_{|V|}}|^{2}$ gives the probability to find the many-body system in excitation configuration |$\ket{n_{1},\ldots,n_{|V|}}$. If we start from a state |$\ket{n_{1},\ldots,n_{|V|}}$, under our assumptions, the next state is essentially $H\ket{n_{1},\ldots,n_{|V|}}$. Therefore, the transition probabilities are essentially the (modulus square) of the entries of $H$.
@@
The last ingredient in our inductive bias is the observation that the operators $h(j_{1},\ldots,j_{i},k_{1},\ldots,k_{o})a^{\dagger}_{k_{1}}\cdots a^{\dagger}_{k_{o}}\cdot a_{j_{1}}\cdots a_{j_{i}}$ also act on excitation configurations |$\ket{n_{1},\ldots,n_{|V|}}$ that have excitations $n_{x}=1$ for $x\neq j_{1},\ldots,j_{i},k_{1},\ldots,k_{o}$, but leave $n_{x}$ unchanged. Similarly, we also apply $h$-values $h_{(i,o)}(\{c_{j_{1}},\ldots,c_{j_{i}}\},\{c_{k_{1}},\ldots,c_{k_{o}}\})$ to excitation configurations that have additional excitations in unrelated atomic clips $c_{x}$.
VI.2 First Steps Towards Quantum mePS
@@
As we described in Section VI.1, the dynamics of such quantum systems are described by time evolution operators of the form $e^{-itH}$ (or more generally, $\mathcal{T}e^{-i\int_{t_{1}}^{t_{2}}H(t)\mathrm{d}t}$, where $\mathcal{T}$ is the time-ordering operator).
We emphasize that most of the considerations in Section VI.1 served to formally develop a classical and discrete analogue. On quantum simulation hardware, we can directly define a parametrized Hamiltonian $H$ of the form in Eq. (8) and define the deliberation as its time evolution, which is a continuous time quantum walk.
@@ VII Discussion and Outlook
The explainability of mePS and the power of our inductive bias were demonstrated in three synthetic environments: two extensions of the Invasion Game and a broken computer diagnosis and repair scenario. The Invasion Game modifications were chosen to visualize in a clean and simple setting the impact of an appropriate many-body inductive bias on the learning process. Using less than necessary excitations leads to bad returns (“underfitting”), while using extra unnecessary excitations slows down the training. In the Computer Maintenance setting, we used a multi-layered agent to provide chains-of-thought of length 2, where we distinguished between the agent’s belief about the causes of problems and the fixes necessary to solve those problems. Hypothesis and plausibility rewards were introduced and applied to different segments of the deliberation path to overcome the credit assignment problem, which tends to occur when blindly applying the basic PS update rule to intermediate layers. The structure of the plausibility reward in particular, encoded causal elements that offered a mechanism for the agent to weakly generalize to unseen percepts. The multi-layer architecture combined with the reward structure helped demonstrate the ease with which a mePS agent can successfully navigate a complex, real-world-inspired environment while maintaining explainability. The inductive bias also proved useful in greatly cutting down the total number of steps and model trainable parameters required to reach the optimal policy.
@@
Acknowledgements.
This research was funded in part by the Austrian Science Fund (FWF) [SFB BeyondC F7102, 10.55776/F71]. For open access purposes, the authors have applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission. We gratefully acknowledge support from the European Union (ERC Advanced Grant, QuantAI, No. 101055129). The views and opinions expressed in this article are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council - neither the European Union nor the granting authority can be held responsible for them.
@@ Appendix A Additional Computer Maintenance Training Details
The external reward shaping function $f(R;b)$ that is used in Subsection V.3 has the form
@@ Appendix B Modifications of the Inductive Bias
In Appendix C, we will show that all inductive biases, including 2SF and 2DP, satisfy Definitions 2 and 5 when the standard probability assignment is used.
In Appendix D, we will prove the promised upper bounds.
@@
    $h_{(2,2)}(\{c_{m},c_{n}\},\{c^{\prime}_{j},c^{\prime}_{k}\})$ such that $j,k\in\{1,2,3,4\}$, $j<k$ and $m,n\in\{1,2,3\}$, $m<n$
@@
    $h_{(2,2)}(\{c_{m},c_{n}\},\{c_{j},c^{\prime}_{k}\})$ such that $j,k\in\{1,2,3,4\}$ and $m,n\in\{1,2,3\}$, $m<n$
@@ Appendix D Detailed Complexity Results and Proofs
Now, assume that the claim is true for all layered many-body hypergraphs that have up to $D-1$ layers and that we have $D$ layers.
@@
By definition of Inductive Bias 2FF (modifying Inductive Bias 1) all deliberations end when all excitations are in layer $L_{D}$, or earlier. The number of excitations cannot increase. Each time a transition is sampled, at least one excitation is removed or moved to the next layer. Moving an excitation to the final layer takes at most $D-1$ steps. There are at most $\sum_{j=1}^{D}|L_{j}|$ excitations, each of which requires no more than $D-1$ steps to be removed or moved to the final layer. In total, we require no more than $(D-1)\times\sum_{j=1}^{D}|L_{j}|$-steps. ∎
@@
Consider a layered many-body mePS agent with layers $(L_{1},\dots L_{D})$ obeying Inductive Bias 2DP. Then the maximal number of deliberation steps is upper-bounded by $(D-1)$.

### v4

@@ 7.1 Full Quantum Motivation
[…]
@@
$H=\sum_{o,i}\sum_{j_{1},\ldots,j_{i}}\sum_{k_{1},\ldots,k_{o}}$ $h(j_{1},\ldots,j_{i},k_{1},\ldots,k_{o})\,a^{\dagger}_{k_{1}}\ldots a^{\dagger}_{k_{o}}\cdot a_{j_{1}}\ldots a_{j_{i}}$ (10)
@@
For our many-body physics-inspired inductive bias, we identify each term $|h(j_{1},\ldots,j_{i},k_{1},\ldots,k_{o})|^{2}$ from $H$ with an unnormalized transition probability/$h$-value $h_{(i,o)}(\{c_{j_{1}},\ldots,c_{j_{i}}\},\{c_{k_{1}},\ldots,c_{k_{o}}\})\in\mathbb{R}$. Here, the set symbol $\{\}$ indicates that the order of atomic clips does not matter (in many-body physics, all $a_{j}$ commute or anti-commute with each other). A crucial part of our inductive bias is that we demand a cutoff for $i$ and $o$. One can also introduce further physics-inspired assumptions such as particle number conservation, formalized as $i=o$.
[…]
7.2 First Steps Towards Quantum MEPS
@@
As we described in Section 7.1, the dynamics of such quantum systems are described by time evolution operators of the form $e^{-itH}$ (or more generally, $\mathcal{T}e^{-i\int_{t_{1}}^{t_{2}}H(t)\mathrm{d}t}$, where $\mathcal{T}$ is the time-ordering operator).
We emphasize that most of the considerations in Section 7.1 served to formally develop a classical and discrete analogue. On quantum simulation hardware, we can directly define a parametrized Hamiltonian $H$ of the form in Eq. (10) and define the deliberation as its time evolution, which is a continuous time quantum walk.
@@ 8 Discussion and Outlook
The explainability of MEPS and the power of our inductive bias were demonstrated in three synthetic environments: two extensions of the Invasion Game and a broken computer diagnosis and repair scenario. The Invasion Game modifications were chosen to visualize in a clean and simple setting the impact of an appropriate many-body inductive bias on the learning process. Using less than necessary excitations leads to bad returns (“underfitting”), while using extra unnecessary excitations slows down the training. In the Computer Maintenance setting, we used a multi-layered agent to provide chains-of-thought of length 2, where we distinguished between the agent’s belief about the causes of problems and the fixes necessary to solve those problems. Hypothesis and plausibility rewards were introduced and applied to different segments of the deliberation path to overcome the credit assignment problem, which tends to occur when blindly applying the basic PS update rule to intermediate layers. The structure of the plausibility reward in particular, encoded causal elements that offered a mechanism for the agent to weakly generalize to unseen percepts. The multi-layer architecture combined with the reward structure helped demonstrate the ease with which a MEPS agent can successfully navigate a complex, real-world-inspired environment while maintaining explainability. The inductive bias also proved useful in greatly cutting down the total number of steps and model trainable parameters required to reach the optimal policy, in particular, compared to non-PS algorithms such as Q-learning and its tabular PS-inspired multi-layer extension introduced in this work.
@@
9 Acknowledgements
This research was funded in part by the Austrian Science Fund (FWF) [10.55776/F7102, 10.55776/WIT9503323]. For open access purposes, the authors have applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission. We gratefully acknowledge support from the European Union (ERC Advanced Grant, QuantAI, No. 101055129). The views and opinions expressed in this article are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council - neither the European Union nor the granting authority can be held responsible for them.
@@ Appendix A Additional Computer Maintenance Training Details
The external reward shaping function $f(R;b)$ that is used in Subsection 6.3 has the form
@@ Appendix B Modifications of the Inductive Bias
In  C, we will show that all inductive biases, including 2SF and 2DP, satisfy Definitions 2 and 5 when the standard probability assignment is used.
In  D, we will prove the promised upper bounds.
@@
    $h_{(2,2)}(\{c_{m},c_{n}\},\{c^{\prime}_{j},c^{\prime}_{k}\})$ such that $\in\{1,2,3,4\}, j<kj<k$ and $\in\{1,2,3\}, m<nm<n$
@@
    $h_{(2,2)}(\{c_{m},c_{n}\},\{c_{j},c^{\prime}_{k}\})$ such that $j,k\in\{1,2,3,4\}$ and $\in\{1,2,3\}, m<nm<n$
@@ Appendix D Detailed Complexity Results and Proofs
Now, assume that the claim is true for all layered many-body hypergraphs that have up to D1D-1 layers and that we have $D$ layers.
@@
By definition of Inductive Bias 2FF (modifying Inductive Bias 1) all deliberations end when all excitations are in layer $L_{D}$, or earlier. The number of excitations cannot increase. Each time a transition is sampled, at least one excitation is removed or moved to the next layer. Moving an excitation to the final layer takes at most D1D-1 steps. There are at most $\sum_{j=1}^{D}|L_{j}|$ excitations, each of which requires no more than D1D-1 steps to be removed or moved to the final layer. In total, we require no more than $(D-1)\times\sum_{j=1}^{D}|L_{j}|$-steps. ∎
@@
Consider a layered many-body MEPS agent with layers $(L_{1},\dots L_{D})$ obeying Inductive Bias 2DP. Then the maximal number of deliberation steps is upper-bounded by (D1)(D-1).